OpenStudy (anonymous):

I need help on this question. Lesley's trash can had a diameter of 18.25 inches. What was the circumference of the can?

OpenStudy (anonymous):

1. Take half of 18.25 2. Then use 9.125 and put it in the circumference formula.

OpenStudy (espex):

\(C = 2*\pi * r\), it also equals \(C = \pi d\)

OpenStudy (anonymous):

Plug in the diameter in the formula \[C =\pi \times d\]

OpenStudy (anonymous):

So it will be:\[C=\pi \times 18.25\]

OpenStudy (anonymous):

What did you get @pattyworth ?